# Panoramic Distortion-Aware Tokenization for Person Detection and Localization Using Transformers in Overhead Fisheye Images

### v2

@@ 3.2 Panoramic distortion-aware tokenization
Figure 2: Panoramic distortion-aware tokenization. The tiling examples indicate tiling failure with gaps and successful tiling in (a) and (b), respectively. The blue lines through the tile centers, where $u=v/3$, indicate that the heights of the tiles change linearly. (c) Our tiling for the tokenization method fills a feature map. (d) The unit figure of the self-similarity figure divisions is indicated when the number of regions $K$ is five.
The goal of PDAT is to match the person’s height with the height of the square tiles that group the tokens using nonuniformity divisions. As described in Section 3.1, the person’s height decreases linearly at incident angles exceeding 70, corresponding to approximately the top one-fourth of the height of the equirectangular images. Therefore, we align the square tiles, which have sizes that depend on the vertical axis $v$. We found that use of inappropriate tile sizes leads to gaps and use of consecutive half-sizes fills the feature maps, as shown in Figure 2(a) and (b), respectively. Our overall tiling method is illustrated in Figure 2(c). The top one-fourth of the image is occupied by performing consecutive half-size tiling, and the other three-fourths consist of tiles of the same size. Each of these tiles has $a_{k}\times a_{k}$ feature vectors, where $a_{k}$ is the side length of the number of feature vectors in a single tile of the $k$-th region, as shown in Figure 2(d). Our tiling is defined on the basis of the self-similarity figures when using $K$ regions, with the exception that the top two rows have tiles of the same size because we have used finite divisions, as illustrated in Figure 2(d). The unit figure consists of a square and two rectangles, and each rectangle can be divided recursively using another set composed of a square and two rectangles. Note that our division method was inspired by the Cantor ternary set [46], but our figure is not a fractal according to Mandelbrot’s definition [31]. Therefore, our tokenization method using self-similarity figures represents the optimal division in terms of matching the person’s height with the tile height.
@@ 3.3 Network architecture
Panoramic image conversion. Our method converts an input fisheye image into an equirectangular panoramic image via remapping, as illustrated in Figure 1. When the input image is rectangular, we center-crop the image. To address any uncalibrated cameras, we assume that the input images represent the stereographic projection, which is a standard fisheye camera model. In the case of radial distortion, this projection uses a trigonometric function: $\gamma=2f_{w}\cdot\tan(\theta/2)$, where $f_{w}$ is the focal length in Equation (4) and $\theta$ represents the incident angle in Section 2 (visual localization). We can then determine the focal length by fitting the image circle using $90^{\circ}$ incident angles. The image coordinates $\tilde{\mathbf{u}}$ in the fisheye images are projected onto a unit sphere in world coordinates $\tilde{\mathbf{p}}$ using backprojection [54], and then the world coordinates $\tilde{\mathbf{p}}$ are projected onto an equirectangular image with an arbitrary image width [42]. We use panoramic images with incident angles from $0^{\circ}$ to $90^{\circ}$ because the incident angle range from $90^{\circ}$ to $180^{\circ}$ represents the directions behind the cameras.
Transformer-based detector.
@@
Figure 3: Overall proposed network architecture. (a) The network architecture has a detector head and a backbone consisting of four stages. The parameter $N$ indicates the number of blocks in the subscript stage. (b) The panoramic distortion-aware tokenization process aggregates the significance map $\mathbf{S}$ and the feature map $\mathbf{X}$ using max pooling.
@@
Following the approach in [64], we use the dynamic anchor box detection with transformers (DAB-DETR) [27], which has shown strong object detection performances. Window-based transformers are not suitable for use in the backbone of DAT-DETR because PDAT divides the feature maps nonuniformly. Therefore, we use an aggregation-based transformer called SG-Former [37] for the backbone. Note that SG-Former is the only aggregation-based method of this type available. The proposed network architecture is shown in Figure 3(a). The detector head block represents DAB-DETR without the backbone, and the remaining blocks represent SG-Former and PDAT. The patch embedding is extracted using the CNN, and the transformer block is a vanilla transformer block [9]. The DAB-DETR method and the SG-Former follow the details given in the corresponding papers in the literature. We explain integration of PDAT into SG-Former in the following two steps: 1) SG-Former attention and 2) our tokenization details.
@@
SG-Former attention. The core ideas of the SG-Former are hybrid-scale self-attention and self-guided attention. The hybrid-scale attention process extracts both global and local information from a feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ are the height, the width, and the channels of the map, respectively. To obtain the local information, down-sampling is applied after the self-guided transformer blocks. The hybrid-scale attention process also calculates the significance map $\mathbf{S}\in\mathbb{R}^{H\times W}$, based on the feature map. For example, the face is a high-significance area and the background is a low-significance area. The self-guided attention process aims to aggregate the feature vectors based on the significance map. For example, the feature vectors in a high-significance area of the face are aggregated into half the number of feature vectors; in contrast, the feature vectors in a low-significance area of the background are aggregated into 1/28th feature vectors.
@@
Our tokenization details. We aim to extend the self-guided attention process using PDAT. Our modification of SG-Former inserts PDAT between the hybrid-scale transformer block and the self-guided block. The PDAT method then leverages the significance map $\mathbf{S}$ based on the feature vectors grouped by the tiles. This approach extends the self-guided attention procedure seamlessly. Figure 3(b) shows the PDAT pipeline. First, the significance maps are divided using our tiling method, as described in Section 3.2(c). From top to bottom, we split these maps horizontally into strips called regions. Second, we extract 2D max pooling indices for each region $k$ using square kernels of $a_{k}\times a_{k}$ in each tile; i.e., the kernel size is dependent on the region. Finally, we leverage the significance values extracted by max pooling. This leveraging in this case is that the significance values are multiplied by the scaling factor $\alpha$. This leveraging approach is designed to preserve the significance values of smaller people during aggregation of the feature vectors because larger people in the images yield multiple areas with high significance values. Therefore, our tokenization approach allows us to balance the significance values among people of various sizes.
@@ 3.4 Training and inference
Training. We generate equirectangular images from each training set. In this generation procedure, we avoid splitting the bounding boxes in the vertical image edges by changing the definition of the 0 pan angles. We then train our network using the panoramic images by following the training procedure given in [27].
Inference. First, we convert an input fisheye image into a panoramic image by a remapping process, as illustrated in Figure 3(a). Second, the panoramic image obtained is then fed into our network. Finally, we project the detected bounding boxes onto the fisheye image, on which the shape of each bounding box is a rotated trapezoid. This projection represents the backprojection from a fisheye image to a panoramic image, as described in Section 3.3.
@@ 4 Experiments
To demonstrate the validity and effectiveness of our approach, we conducted extensive experiments using three large-scale datasets [10, 48, 64], which are composed of overhead fisheye images. We used 1024$\times$1024 input resolution by following the settings given in [10, 22, 41, 47, 64]. For evaluation, we converted the trapezoid bounding boxes into rectangles because no other methods listed in Table 1 produced trapezoid outputs.
@@
Method mAP AP50 AP75 APn APm APf APseen APunseen
val
Seidel et al. [41] VISAPP’19 21.8 59.8 07.6 32.8 28.5 02.3 23.2 19.5
Li et al. [22] AVSS’19 28.5 63.3 20.1 46.8 24.2 01.3 33.8 29.3
Tamura et al. [47] WACV’19 34.8 72.1 27.7 51.7 38.8 08.7 38.8 32.9
RAPiD [10]   CVPRW’20 40.3 77.9 34.8 55.3 41.9 09.2 44.7 37.6
OARPD [34] MTA’24 45.5 78.8 46.3 - - - - -
Yang et al. [64] ICCV’23 47.4 82.6 48.4 64.1 54.3 14.1 50.8 45.7
Ours 52.4 84.6 57.6 66.3 57.9 42.9 56.3 50.7
test
Seidel et al. [41] VISAPP’19 20.2 58.2 07.1 32.2 28.3 03.9 22.4 18.9
Li et al. [22] AVSS’19 27.2 65.2 21.3 47.2 24.8 01.3 31.8 27.6
Tamura et al. [47] WACV’19 34.2 72.8 28.7 53.7 37.3 06.5 39.5 33.2
RAPiD [10]   CVPRW’20 39.2 77.9 35.4 54.8 40.1 07.9 44.2 37.3
OARPD [34] MTA’24 45.8 79.4 45.8 - - - - -
Yang et al. [64] ICCV’23 46.2 81.1 47.3 66.1 53.5 12.6 49.3 44.9
Ours 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
4.1 Parameter settings
For our network backbone, we used SG-Former-S [37], which has the smallest backbone size among the S, M, and B types, and was pretrained on ImageNet [38]. We trained our network using a combination of a mini-batch size of 8 for 120 epochs and the AdamW optimizer [30]. The initial learning rate was set at $2\times 10^{-4}$ and was multiplied by 0.1 at the 40th epoch. We used the hyperparameters $K=5$ and $\alpha=2$. Because the majority of bounding boxes were near the image circles of fisheye images, we used the 3072-pixel width of panoramic images, approximately corresponding to the circumference of fisheye images with 1024 $\times$ 1024 pixels: $\pi\times 1024=3217\approx 3072$ pixels. Unless otherwise specified, we used panoramic images with 3072 $\times$ 768 pixels.
4.2 Experimental results
We used the person detection and localization results reported by Yang et al. [64] in the methods of Seidel et al. [41], Li et al. [22], Tamura et al. [47], RAPiD [10], and Yang et al. [64]. We also used the OARPD results reported in [34]. Unless otherwise specified, we refer to these results throughout. Because of the lack of previous performance reports, we implemented methods for comparison of Tamura et al. [47] and OARPD [34], based on the corresponding papers using PyTorch [32]. In addition, the official codes from Yang et al. [64] and RAPiD [10] [^1] Yang et al.: https://github.com/BUPT-PRIV/LOAF,  RAPiD: https://github.com/duanzhiihao/RAPiD were used for evaluation.
4.2.1 Person detection on LOAF
@@
Seidel et al. [41] 20.9 50.6 10.2 80.6 39.5 53.0 16.1 39.4 09.0 70.9 38.6 50.0
Li et al. [22] 34.2 75.7 28.6 86.3 65.4 74.4 25.2 69.9 30.2 81.4 64.5 72.0
Tamura et al. [47] 1 28.8 61.6 23.0 73.1 65.8 69.3 12.8 27.2 10.9 53.0 34.3 41.6
RAPiD [10] 1 38.3 78.6 30.8 91.8 80.3 85.7 25.0 54.0 17.4 83.7 57.6 68.2
OARPD [34] 1 11.4 28.9 06.6 80.5 31.3 45.1 04.1 10.0 02.2 28.8 19.0 22.9
Yang et al. [64] 1 39.7 78.3 37.1 82.9 72.6 77.4 20.5 45.3 16.3 66.8 43.0 52.3
@@
Baseline (DAB-DETR) 40.2 74.0 39.6 48.2 49.5 32.7 42.5 38.7
$+$ Panoramic conversion 49.8 84.7 52.7 59.3 55.3 39.7 51.0 48.9
$+$ PDAT 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
Table 5: Panoramic-image resolution analysis on the LOAF test set
 Panoramic image width mAP AP50 AP75 APn APm APf APseen APunseen
@@
Metrics and protocols. To demonstrate the validity and the effectiveness of the proposed method on LOAF, we used the mean average precision (mAP) following a COCO evaluation [24]. We used the subscripts $n$, $m$, and $f$ to denote the near (0–10 m), mid-range (10–20 m), and distant (over 20 m) person positions, respectively, e.g., APn. LOAF also provides seen and unseen scenes for test splits using the subscripts $seen$ and $unseen$, respectively. We used the official splits of the training, validation, and test sets.
Results. Our method achieved the highest mAP among the methods listed in Table 2 when used on both the validation and test sets. The mAP obtained using our method was substantially higher than that obtained using Yang et al.’s method [64] by values of 5.0 and 8.3 on the LOAF validation and test sets, respectively. In addition, the APf value obtained using our method on the test set was notably larger than that of Yang et al.’s method [64] by 32.0, where 67% of the bounding boxes were labeled as $f$: over 20 m. Our method also achieved an APseen of 56.2 and an APunseen of 53.6 on the test set, thus outperforming the other methods. The substantial margin between our method and Yang et al.’s method [64] on the unseen test set of 8.7 demonstrates the scene robustness.

### v3

@@ 3.3 Panoramic distortion-aware tokenization
On the basis of the design discussed above, we propose PDAT that addresses person detection in overhead fisheye images. The key insight of PDAT is that it maintains consistent person sizes relative to tile sizes across different regions, regardless of their incident angles. To achieve this goal, PDAT employs nonuniform divisions that adapt to the varying sizes of persons at different incident angles, as shown in Figure 3(a). These tiles are designed to fill a panoramic feature map without gaps, ensuring complete coverage for reliable person detection.
As shown in Figure 2(b), the person’s height exhibits an approximately linear decrease at incident angles exceeding 70. This portion corresponds to about the top quarter of the height of the equirectangular images. On the basis of this observation, we align square tiles whose sizes depend on the vertical axis. We found that use of consecutive half-sizes fills the feature maps, as illustrated in Figure 3(a). The top quarter of the image is occupied by performing consecutive half-size tiling, and the other three-quarters consist of tiles of uniform size because the bounding box heights show nonlinear trends in this region with fewer instances.
Each of these square tiles contains $a_{k}\times a_{k}$ feature vectors, where $a_{k}$ represents the number of feature vectors along each side of a tile in the $k$th region, as shown in Figure 3(b). Our tiling is defined on the basis of the self-similar figures when using $K$ regions, with the exception that the top two rows in region 1 have tiles of the same size due to finite divisions. The unit figure consists of a square and two rectangles, and each rectangle can be divided recursively using another set composed of a square and two rectangles, as illustrated in the rightmost part of Figure 3(b). Note that our division method was inspired by the Cantor ternary set [Smith_LMS_1874], but our self-similar figure is not a fractal according to Mandelbrot’s definition [Mandelbrot_Science_1967].
Although the tile sizes vary among regions 3, 4, and 5, as shown in Figure 3(a), the person-bounding boxes appear similar in size relative to their respective tiles. This design maintains similar complexity across all regions, despite the varying absolute sizes of persons in the panoramic image. As illustrated in Figure 3(b), our tokenization method uses recursive divisions to maintain consistent person-to-tile size ratios, particularly in portions with large incident angles.
@@ 3.4 Network architecture
Figure 4: Overall network architecture of our proposed method. The model consists of a detector head and a backbone with four stages. The NN in the subscript stage indicates the block count. The $\mathbf{S}$ and $\mathbf{X}$ are the significance map and feature map, respectively.
Our network architecture consists of four main components: panoramic image conversion, a detector head based on the dynamic anchor box detection with transformers (DAB-DETR) [Liu_ICLR_2022], a backbone based on SG-Former [Ren_ICCV_2023], and our proposed PDAT.
Panoramic image conversion. Our method converts an input fisheye image into an equirectangular panoramic image via remapping, as illustrated in Figure 4. When the input image is rectangular, we center-crop the image. To address any uncalibrated cameras, we assume that the input images represent the stereographic projection, which is a standard fisheye camera model. In the case of radial distortion [Wakai_ECCV_2022], this projection uses a trigonometric function: $\gamma=2f_{w}\cdot\tan(\theta/2)$, where $\gamma$ is the distortion, $f_{w}$ is the focal length, and $\theta$ represents the incident angle, as shown in Figure 2(a). We can then determine the focal length by fitting the image circle using $90^{\circ}$ incident angles. The image coordinates in the fisheye images are projected onto a unit sphere in world coordinates using backprojection [Wakai_ECCV_2022], and then the world coordinates are projected onto an equirectangular image with an arbitrary image width [Sharma_AIVR_2019]. We use panoramic images with incident angles from $0^{\circ}$ to $90^{\circ}$ because the incident angle range from $90^{\circ}$ to $180^{\circ}$ represents the directions behind the cameras.
@@
Detector and backbone selection. Following the approach in [Yang_ICCV_2023], we use DAB-DETR [Liu_ICLR_2022], which has shown strong object detection performances. Window-based transformers are not suitable for use in the backbone of DAB-DETR because PDAT divides the feature maps nonuniformly. Therefore, we use an aggregation-based transformer called SG-Former [Ren_ICCV_2023] for the backbone. Note that SG-Former is the only aggregation-based method of this type available. The proposed network architecture is shown in Figure 4. The detector head block represents DAB-DETR without the backbone, and the remaining blocks represent SG-Former and PDAT. The patch embedding is extracted using the CNN, and the transformer block is a vanilla transformer block [Dosovitskiy_ICLR_2021]. The implementations of DAB-DETR and SG-Former follow the details provided in their respective papers. We explain the integration of PDAT into SG-Former in the following two steps: 1) SG-Former attention mechanism and 2) PDAT integration details.
@@
SG-Former attention mechanism. The core ideas of the SG-Former are hybrid-scale self-attention and self-guided attention. The hybrid-scale attention process extracts both global and local information from a feature map $\mathbf{X}\in\mathbb{R}^{H\times W\times C}$, where $H$, $W$, and $C$ are the height, the width, and the channels of the map, respectively. The hybrid-scale attention process also calculates the significance map $\mathbf{S}\in\mathbb{R}^{H\times W}$, based on the feature map. For example, the face is a high-significance area and the background is a low-significance area. The self-guided attention process aims to aggregate the feature vectors based on the significance map. For example, the feature vectors in a high-significance area of the face are aggregated into half the number of feature vectors; in contrast, the feature vectors in a low-significance area of the background are aggregated into 1/28th feature vectors. To obtain multi-scale features, down-sampling is applied between stages.
@@
PDAT integration details. We aim to extend the self-guided attention process using PDAT. Our modification of SG-Former inserts PDAT between the hybrid-scale transformer block and the self-guided block, as shown in Figure 4. The PDAT method leverages the significance map $\mathbf{S}$ based on the feature vectors grouped by the tiles. The integration process consists of the following steps. First, the significance maps are divided using our tiling method, as described in Section 3.3. From top to bottom, we split these maps horizontally into strips called regions, as shown in Figure 3(b). Second, we extract 2D max pooling indices for each region $k$ using square kernels of $a_{k}\times a_{k}$ with horizontal strides $a_{k}$ in each tile. The $a_{k}$ corresponds to the height of region $k$ in the upper quarter of the image, with the exception that region 1 uses its half height, while a fixed height of $H/4$ is used for the remaining three-quarters. Finally, we leverage the significance values extracted by the max pooling from the second step. In this leveraging, the significance values are multiplied by the scale factor $\alpha$.
This approach seamlessly integrates into the self-guided attention procedure to preserve the significance values of smaller persons in the feature vector aggregation process. We introduce this scaling because larger persons dominate the attention mechanism by generating multiple high-significance areas in SG-Former without PDAT.
@@ 3.5 Training and inference
Training. We generate equirectangular images from each training set. In this generation procedure, we avoid splitting the bounding boxes in the vertical image edges by changing the definition of the 0 azimuth angles. We then train our network using the panoramic images by following [Liu_ICLR_2022].
Inference. First, we convert an input fisheye image into a panoramic image by a remapping process, as illustrated in Figure 4. Second, the remapped panoramic image is then fed into our network. Finally, we project the detected bounding boxes onto the fisheye image, on which the shape of each bounding box is a rotated trapezoid. This projection represents the backprojection from a fisheye image to a panoramic image.
@@ 4 Experiments
Method val test
mAP AP50 AP75 APn APm APf APseen APunseen mAP AP50 AP75 APn APm APf APseen APunseen
Seidel et al[Seidel_VISAPP_2019] 21.8 59.8 07.6 32.8 28.5 02.3 23.2 19.5 20.2 58.2 07.1 32.2 28.3 03.9 22.4 18.9
Li et al[Li_AVSS_2019] 28.5 63.3 20.1 46.8 24.2 01.3 33.8 29.3 27.2 65.2 21.3 47.2 24.8 01.3 31.8 27.6
Tamura et al[Tamura_WACV_2019] 34.8 72.1 27.7 51.7 38.8 08.7 38.8 32.9 34.2 72.8 28.7 53.7 37.3 06.5 39.5 33.2
RAPiD [Duan_CVPRW_2020] 40.3 77.9 34.8 55.3 41.9 09.2 44.7 37.6 39.2 77.9 35.4 54.8 40.1 07.9 44.2 37.3
OARPD [Qiao_MTA_2024] 45.5 78.8 46.3 - - - - - 45.8 79.4 45.8 - - - - -
Yang et al[Yang_ICCV_2023] 47.4 82.6 48.4 64.1 54.3 14.1 50.8 45.7 46.2 81.1 47.3 66.1 53.5 12.6 49.3 44.9
Ours 52.4 84.6 57.6 66.3 57.9 42.9 56.3 50.7 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
@@
Seidel et al[Seidel_VISAPP_2019] VISAPP’19 20.9 50.6 10.2 80.6 39.5 53.0 16.1 39.4 09.0 70.9 38.6 50.0
Li et al[Li_AVSS_2019] AVSS’19 34.2 75.7 28.6 86.3 65.4 74.4 25.2 69.9 30.2 81.4 64.5 72.0
Tamura et al[Tamura_WACV_2019] 1 WACV’19 28.8 61.6 23.0 73.1 65.8 69.3 12.8 27.2 10.9 53.0 34.3 41.6
RAPiD [Duan_CVPRW_2020] 1 CVPRW’20 38.3 78.6 30.8 91.8 80.3 85.7 25.0 54.0 17.4 83.7 57.6 68.2
OARPD [Qiao_MTA_2024] 1 MTA’24 11.4 28.9 06.6 80.5 31.3 45.1 04.1 10.0 02.2 28.8 19.0 22.9
Yang et al[Yang_ICCV_2023] 1 ICCV’23 39.7 78.3 37.1 82.9 72.6 77.4 20.5 45.3 16.3 66.8 43.0 52.3
@@
To demonstrate the validity and effectiveness of our approach, we conducted extensive experiments using three large-scale datasets [Duan_CVPRW_2020, Tezcan_WACV_2022, Yang_ICCV_2023], which are composed of overhead fisheye images. We used $\times1024$ input resolution by following the settings given in [Duan_CVPRW_2020, Li_AVSS_2019, Seidel_VISAPP_2019, Tamura_WACV_2019, Yang_ICCV_2023]. For evaluation, we converted the trapezoid bounding boxes into rectangles because no other methods listed in Table 1 produced trapezoid outputs.
4.1 Parameter settings
[…]
4.2 Person localization
For person localization, we adopt the geometry-based method proposed by Yang et al[Yang_ICCV_2023]. This method estimates person locations by calculating the intersection point between the incident ray from the camera and the ground plane, where the incident ray is determined from the center of the lower edge of a detected person-bounding box.
4.3 Experimental results
We used the person detection and localization results reported by Yang et al[Yang_ICCV_2023] in the methods of Seidel et al[Seidel_VISAPP_2019], Li et al[Li_AVSS_2019], Tamura et al[Tamura_WACV_2019], RAPiD [Duan_CVPRW_2020], and Yang et al[Yang_ICCV_2023]. We also used the OARPD [Qiao_MTA_2024] results reported in [Qiao_MTA_2024]. Unless otherwise specified, we refer to these results throughout. Because of the lack of previous performance reports, we implemented methods for comparison of Tamura et al[Tamura_WACV_2019] and OARPD [Qiao_MTA_2024], based on the corresponding papers using PyTorch [Paszke_NIPS_2019]. In addition, the official codes from Yang et al[Yang_ICCV_2023] and RAPiD [Duan_CVPRW_2020] were used for evaluation.
4.3.1 Person detection on LOAF
@@
  ++ Baseline 1 40.2 74.0 39.6 48.2 49.5 32.7 42.5 38.7
++ Panorama 2 49.8 84.7 52.7 59.3 55.3 39.7 51.0 48.9
++ PDAT 54.5 87.2 61.1 62.8 60.5 44.6 56.2 53.6
1 “Baseline” denotes DAB-DETR using fisheye images.
2 “Panorama” denotes panoramic conversion.
Table 5: Panoramic-image width analysis on the LOAF test set
  Image width mAP AP50 AP75 APn APm APf APseen APunseen
@@
Metrics and protocols. To demonstrate the validity and the effectiveness of the proposed method on LOAF, we used the mean average precision (mAP) following the COCO evaluation [Lin_ECCV_2014]. We used the subscripts $n$, $m$, and $f$ to denote the near (0–10 m), mid-range (10–20 m), and distant (over 20 m) person positions, respectively, e.g., APn. LOAF also provides seen and unseen scenes for test splits using the subscripts seenseen and unseenunseen, respectively. We used the official splits of the training, validation (val), and test sets.
Results. Our method achieved the highest mAP among the methods listed in Table 2 when used on both the validation and test sets. The mAP obtained using our method was substantially higher than that obtained using Yang et al.’s method [Yang_ICCV_2023] by values of 5.0 and 8.3 on the LOAF validation and test sets, respectively. In addition, the APf value obtained using our method on the test set was notably larger than that of Yang et al.’s method [Yang_ICCV_2023] by 32.0, where 67% of the bounding boxes were labeled as $f$: over 20 m. Our method also achieved an APseen of 56.2 and an APunseen of 53.6 on the test set, thus outperforming the other methods. The substantial margin between our method and Yang et al.’s method [Yang_ICCV_2023] on the unseen test set of 8.7 demonstrates the scene robustness.